# FIDLAR: Forecast-Informed Deep Learning Architecture for Flood Mitigation

### v1

@@ Abstract
In coastal river systems, frequent floods, often occurring during major storms or king tides, pose a severe threat to lives and property. However, these floods can be mitigated or even prevented by strategically releasing water before extreme weather events with hydraulic structures such as dams, gates, pumps, and reservoirs. A standard approach used by local water management agencies is the “rule-based” method, which specifies predetermined pre-releases of water based on historical and time-tested human experience, but which tends to result in excess or inadequate water release. The model predictive control (MPC), a physics-based model for prediction, is an alternative approach, albeit involving computationally intensive calculations. In this paper, we propose a Forecast Informed Deep Learning Architecture, FIDLAr, to achieve rapid and optimal flood management with precise water pre-releases. FIDLAr seamlessly integrates two neural network modules: one called the Flood Manager, which is responsible for generating water pre-release schedules, and another called the Flood Evaluator, which assesses these generated schedules. The Evaluator module is pre-trained separately, and its gradient-based feedback is used to train the Manager model, ensuring optimal water pre-releases. We have conducted experiments using FIDLAr with data from a flood-prone coastal area in South Florida, particularly susceptible to frequent storms. Results show that FIDLAr is several orders of magnitude faster than currently used physics-based approaches while outperforming baseline methods with improved water pre-release schedules. Our code is at https://github.com/JimengShi/FIDLAR/.
@@ 1 Introduction
Floods can result in catastrophic consequences with considerable loss of life Jonkman and Vrijling (2008), huge socio-economic impact Wu et al. (2021), property damage Brody et al. (2007), and environmental devastation Yin et al. (2023a). They pose a threat to food and water security as well as sustainable development Kabir and Hossen (2019). What is even more alarming is that research indicates global climate change may lead to a drastic increase in flood risks in terms of both frequency and scale Wing et al. (2022); Hirabayashi et al. (2013). In coastal areas, flood risks are increasing due to sea-level rise Sadler et al. (2020). Therefore, effective flood management is of utmost importance.
@@
To mitigate flood risks, water management agencies have built controllable hydraulic structures such as dams, gates, pumps, and reservoirs in river systems Kerkez et al. (2016). Determining optimal control schedules of these hydraulic structures is however a challenging problem Bowes et al. (2021). Rule-based methods Bowes et al. (2021); Sadler et al. (2019) formulate control schedules based on insights gained from historically observed data. These schedules are often formalized and extensively documented in databases Marchese et al. (2018). They represent the collective wisdom gathered over decades of experience in managing specific river systems and are commonly adopted by local water management authorities. However, these schedules may expose vulnerabilities during extreme events, which are becoming more likely due to climate change, and might lack sufficient locally available historical data, particularly for events like 500- or 1000-year floods. Moreover, these schedules may not offer effective solutions for complex river systems Schwanenberg et al. (2015).
Model Predictive Control (MPC) is a well-known strategy that tackles this challenge as an optimization process Karimanzira (2016). By optimizing variables such as water levels and discharges, the algorithm generates control schedules for hydraulic structures to effectively mitigate overall flood risks Delaney et al. (2020); Zarei et al. (2021). Random initialization followed by soft-computing techniques such as genetic algorithms is often employed to perform the optimizations. Subsequently, physics-based models such as HEC-RAS and SWMM are used to assess the generated control schedules. Publications on this topic include the following Leon et al. (2020); Sadler et al. (2019); Vermuyten et al. (2018); Chen et al. (2016); Leon et al. (2014); Yin et al. (2023b). However, such MPC methods are prohibitively slow since they require thousands of simulations using physics-based models.
In the last decade or so, Machine learning (ML) has emerged as a powerful tool in this domain. So far, the ML-based methods have been applied to flood prediction Mosavi et al. (2018); Shi et al. (2023b), flood detection Tanim et al. (2022); Shahabi et al. (2020), susceptibility assessment Saha et al. (2021); Islam et al. (2021), and post-flood management Munawar et al. (2019). However, ML-based methods have not been applied to flood mitigation tasks. We aim to fill this gap. Moreover, we replace the compute-intensive MPC approach with a framework based purely on deep learning (DL), where both the generator and evaluator are achieved by DL modules.
Gradient-based planning with world models aims to seek optimal actions by training a world model and conducting inferences using that trained model. We refer the reader to Figure 1(a) from Jyothir et al. (2023).
@@
Figure 1: Forecast-Informed Deep Learning Architecture (FIDLAr). Input data in the form of a table consisting of five variables as shown in the left end of the diagram. The variables $w$ and $k$ are the lengths of the past and prediction windows, respectively. The parts of the table colored green are provided as inputs to the programs, while the orange and blue parts (with question marks) are outputs. The black and red arrows represent the propagation and back-propagation processes, respectively. The Flood Manager and Flood Evaluator represent DL models, the former to predict the control schedules of gates and pumps to pre-release water, and the latter to determine water levels for those control schedules. Loss functions, $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$, penalize the flooding and water wastage beyond pre-specified thresholds, respectively.
Our proposed Forecast Informed Deep Learning Architecture, FIDLAr, uses gradient-based planning to achieve optimal flood management with rapid and feasible water pre-releasing. See Figure 1 for details on its architecture. FIDLAr’s characteristics are summarized as follows:
@@
  • FIDLAr seamlessly combines two DL models in series: Flood Manager and Flood Evaluator. The former model is responsible for generating water pre-release schedules, while the latter model accurately forecasts the resulting water levels. The latter model serves as an evaluator for the output of the former. During training, the gradient feedback from the Evaluator model can be back-propagated to reinforce the Manager model to generate effective water pre-release schedules. The Manager and Evaluator play the role of Planner and World Model, respectively, from the model from Jyothir et al. (2023).
@@
  • FIDLAr is a data-driven approach, learning flood mitigation strategies from historically observed data. Once trained, it offers rapid response capabilities. This highlights the advantages of DL-based models over physics-based models, particularly in the context of real-time flood management.
@@
  • FIDLAr is a model-agnostic framework, where both the Manager and Evaluator could be any type of DL model with differentiable loss functions that allow back-propagation.
@@
  • FIDLAr is trained with a custom loss function to balance both flood risks and water wastage at the same time, which is a multitask learning method.
  • FIDLAr makes effective use of reliable forecasts of specific variables (e.g., precipitation and tidal information) for the near future.
  • The explainability features of FIDLAr allow us to better understand the behavior of complex river systems.
@@ 2 Problem Formulation
Flood mitigation involves the prediction of control schedules for hydraulic structures like gates and pumps within the river system, denoted as $X^{gate,pump}_{t+1:t+k}$, spanning $k$ time points from $t+1$ to $t+k$ into the future. This prediction uses as input historical input data (on all variables), $X$, from the preceding $w$ time points, in conjunction with reliably forecasted covariates (such as rainfall and tide) for the next $k$ time points. Thus, we train a DL model whose transfer function, $\mathcal{F}_{\theta}$, has the form:
$\mathcal{F}_{\theta}:(X_{t-w+1:t},X^{cov}_{t+1:t+k})\rightarrow X^{gate,pump}_{t+1:t+k},$ (1)
@@ 3 Methodology
3.1 The Challenge
Intuitively, an ML model can be trained to learn the function $f_{\theta}$. However, the challenge here is that the historical data may not represent the desired values for the control schedules of the gates and pumps, especially if it resulted in flood conditions, and therefore cannot be used as the ground-truth values used in traditional supervised learning. Therefore, we separate a “planner” (Flood Manager) module from an “evaluator” (Flood Evaluator) module. We first build an independent and accurate evaluator with the historical data, which attempts to learn the consequences of every possible action. We then treat the planner module design as an optimization problem where we optimize a custom loss function, allowing the system to pick the action that would prevent undesirable consequences. Desired states of the river systems avoid floods on one hand while preventing water wastage on the other. Once the Flood Evaluator is trained, it is frozen and used to train the Flood Manager. The evaluator serves to backpropagate the gradient feedback regarding the quality of the schedules generated by the manager, thus guiding the manager to generate more effective schedules of gates and pumps. Such a training process ends up refining the function, $f_{\theta}$, as represented in Eq. (1).
Figure 2: Flood Evaluator. Unlike Figure 1, the parts shaded green (includes historical data and covariates predicted from near future) and orange (representing any control schedule for the gates and pumps) are both used as inputs. Water levels (blue) are the outputs.
@@ 3.2 Flood Evaluator
Flood Evaluator is tasked with accurately forecasting water levels at designated points of interest within river systems for any control schedule of gates and pumps provided to it. See Figure 2. The underlying transfer function of the Evaluator has the form:
$\mathcal{G}_{\theta}:(X_{t-w+1:t},X^{cov}_{t+1:t+k},X^{gate,pump}_{t+1:t+k})\rightarrow X^{water}_{t+1:t+k}.$ (2)
@@
As mentioned earlier, the Evaluator module is trained independently using the historical data to achieve highly accurate water level predictions for any given set of conditions and control schedules. Therefore, once the Evaluator is trained, its parameters are set in stone for the training of the Manager, where it plays the role of a trained “referee”.
@@
Figure 3: Flood Manager. As in Figure 1, the parts shaded green (historical data) are the inputs, and the parts shaded orange are the outputs. The water levels shaded blue are not predicted.
3.3 Flood Manager
As shown in Figure 3, the Flood Manager is trained to produce optimal control schedules for the hydraulic structures (gates and pumps), taking reliably predictable future information (rain, tide) and all historical data. During its training, since no ground truth is available, it is trained using the gradient descent algorithm Ruder (2016); Andrychowicz et al. (2016), using the partial derivatives of the loss function (Eq. (3)), which in turn needs the evaluator.
3.4 Custom Loss Function
Loss functions are critical in steering the learning process. For flood management, since no ground truth is available, independent metrics of performance are needed. An obvious one is the total time (Figure 4 (a)) for which the water levels either exceed the flooding threshold or dip below the water wastage threshold. Another related metric is the extent to which the limits are exceeded to signify the severity of floods or water wastage (Figure 4 (b)). The lower threshold for flood management is important in practice, since it prevents water wastage, thereby supporting irrigation, facilitating navigation, and maintaining ecological balance. It also prevents the optimization methods from trivially recommending the depletion of valuable water resources to prevent future flooding. The above metrics suggest the following components for the loss function for our models. The final loss function is a balanced combination of $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ in Eq. (4).
where $N$ is the number of water level locations of interest; $k$ is the length of prediction horizon; $X^{flood}$ and $X^{waste}$ represent the thresholds for flooding and water wastage; and their capped versions, $\hat{X}^{flood}$ and $\hat{X}^{waste}$ are obtained using the evaluator module. The combined loss function is given by:
$\mathcal{L}_{total}=\alpha\cdot\mathcal{L}_{1}+\beta\cdot\mathcal{L}_{2},$ (4)
where $\alpha/\beta$ dictates the relative importance of $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$.
Figure 4: The two red bars represent an upper threshold of flooding and a lower threshold of wastage. Shown are (a) the time spans when these thresholds are crossed, and (b) the areas between the water level curve and the threshold bars. Violations of the upper (lower, resp.) threshold are captured in $\mathcal{L}_{1}$ ($\mathcal{L}_{2}$, resp.).
Figure 5: Architecture of Graph Transformer Network for Flood Evaluator. Input variables include Rainfall, Pump, Gate, Tide, and Water levels as shown in Figure 6 and Table 1, which are generically denoted by $R,T,G,P$, and $W$. Output is water levels. We use the same architecture for Evaluator and Manager, but with minor changes accordingly of the input and output(see Figs. 2 and 3).
3.5 Training FIDLAr
The entire training process is outlined below. The Evaluator model is trained in a traditional way using the observed historical data as the ground truth for predictions and minimizing a loss function that penalizes deviations from the ground truth. After the Evaluator model is trained, its parameters are frozen and the Manager model is trained. For every control schedule that the Manager generates, $X^{gate,pump}_{t+1:t+k}$, the Evaluator is used to compute the resulting water levels, which is then used to compute the loss function defined in Eq. (4). Gradient descent Ruder (2016) provides the information to be back-propagated to be used to update the parameters of Manager. Thus,
${X^{*}}^{gate,pump}_{t+1:t+k}=\underset{X^{gate,pump}_{t+1:t+k}}{\arg\min}\{\mathcal{L}_{total}\}.\$ (5)
The training details of FIDLAr are given in Algorithm 1.
4 GTN Model
The Manager and Evaluator modules described so far are model agnostic. We tried many existing architectures for them, as discussed in Section 5.2. The Graph Transformer Network (GTN) architecture combines graph neural networks (GNNs), attention-based transformer networks, long short-term memory networks (LSTMs), and the convolutional neural networks (CNNs). GNN and LSTM modules are combined for learning the spatiotemporal dynamics of water levels, while the Transformer and CNN modules focus on extracting feature representations from the covariates. The Attention mechanism is employed to discern the interactions between covariates and water levels, as shown in Eq. (6) below. Figure 5 presents the GTN architecture used (with minor modifications) for the Evaluator and the Manager.
$Atte(Q,K,V){}$ $=softmax(\frac{Q^{cov}(K^{water})^{T}}{\sqrt{d_{k}}})V^{water}$ (6)
$=softmax(\frac{Q^{water}(K^{cov})^{T}}{\sqrt{d_{k}}})V^{cov},$
where $T$ denotes the transpose operation; $water$ and $cov$ represent the variables of water levels and covariates.
@@ 4 GTN Model
Input: recent past data: ${\bf X}_{t-w+1,t}$
@@
Parameter: $\theta_{E},\theta_{M}$: parameters of Evaluator and Manager
Parameter: $w,k:$ length of past and prediction windows
1:  // Train Flood Evaluator, $\mathcal{G}_{\theta_{E}}$
@@
4:     $\text{MiniBatch}\leftarrow(\{{\bf X}_{t-w+1,t},{\bf X}^{cov}_{t+1,t+k}\},{\bf X}^{water}_{t+1,t+k})$
5:     $\hat{{\bf X}}^{water}_{t+1,t+k}\leftarrow\mathcal{G}_{\theta_{E}}({\bf X}_{t-w+1,t},{\bf X}^{cov}_{t+1,t+k})$
@@
10:  return trained Flood Evaluator, $\mathcal{G}_{\theta_{E}}$
11:  // Train Flood Manager, $\mathcal{F}_{\theta_{M}}$, with frozen $\mathcal{G}_{\theta_{E}}$
@@
14:     $\text{MiniBatch}\leftarrow(\{{\bf X}_{t-w+1,t},{\bf X}^{rain,tide}_{t+1,t+k}\},{\bf X}^{gate,pump}_{t+1,t+k})$
15:     $\hat{{\bf X}}^{gate,pump}_{t+1,t+k}\leftarrow\mathcal{F}_{\theta_{M}}({\bf X}_{t-w+1,t},{\bf X}^{rain,tide}_{t+1,t+k})$
16:     $\hat{{\bf X}}^{water}_{t+1,t+k}\leftarrow\mathcal{G}_{\theta_{E}}({\bf X}_{t-w+1,t},{\bf X}^{rain,tide}_{t+1,t+k},\hat{{\bf X}}^{gate,pump}_{t+1,t+k})$
17:     $\mathcal{L}_{E}=\alpha\cdot\mathcal{L}_{1}(\hat{{\bf X}}^{water}_{t+1,t+k})+\beta\cdot\mathcal{L}_{2}(\hat{{\bf X}}^{water}_{t+1,t+k})$
@@
21:  return trained Flood Manager, $\mathcal{F}_{\theta_{M}}$

### v2

@@ Abstract
In coastal river systems, floods, often during major storms or king tides, severely threaten lives and property. However, hydraulic structures such as dams, gates, pumps, and reservoirs exist in these river systems, and these floods can be mitigated or even prevented by strategically releasing water before extreme weather events. A standard approach used by local water management agencies is the “rule-based” method, which specifies predetermined water prereleases based on historical human experience, but which tends to result in excessive or inadequate water release. Iterative optimization methods that rely on detailed physics-based models for prediction are an alternative approach. Whereas, such methods tend to be computationally intensive, requiring hours or even days to solve the problem optimally. In this paper, we propose a Forecast Informed Deep Learning Architecture, FIDLAr, to achieve rapid and near-optimal flood management with precise water prereleases. FIDLAr seamlessly integrates two neural network modules: one called the Flood Manager, which is responsible for generating water pre-release schedules, and another called the Flood Evaluator, which evaluates those generated schedules. The Evaluator module is pre-trained separately, and its gradient-based feedback is utilized to train the Manager model, ensuring near-optimal water pre-releases. We have conducted experiments with a flood-prone coastal area in South Florida. Results show that FIDLAr is several orders of magnitude faster than currently used physics-based approaches while outperforming baseline methods with improved water pre-release schedules.
Codehttps://github.com/JimengShi/FIDLAR
@@ 1 Introduction
Floods can result in catastrophic consequences with considerable loss of life (Jonkman and Vrijling 2008), huge socio-economic impact (Wu et al. 2021), property damage (Brody et al. 2007), and environmental devastation (Yin et al. 2023a). They pose a threat to food and water security as well as sustainable development (Kabir and Hossen 2019). What is even more alarming is that research indicates global climate change may lead to a drastic increase in flood risks in terms of both frequency and scale (Wing et al. 2022), e.g., coastal flood risks due to sea-level rise (Sadler et al. 2020). Thus, effective flood management is of utmost importance.
@@
To mitigate flood risks, water management agencies have built controllable hydraulic structures such as dams, gates, pumps, and reservoirs in river systems (Kerkez et al. 2016). However, determining the optimal control schedules of these hydraulic structures is a challenging problem (Bowes et al. 2021). Rule-based methods (Sadler et al. 2019) formulate control schedules based on insights gained from historically observed data. The rules represent the collective wisdom gathered over decades of experience in managing specific river systems. Nevertheless, these rules may expose vulnerabilities while dealing with extremely rare events and may not offer effective solutions for complex river systems under such conditions (Schwanenberg, Becker, and Xu 2015).
@@
Figure 1: Forecast-Informed Deep Learning Architecture (FIDLAr). Input data consists of five categories of variables as shown in the left table. The variables $w$ and $k$ are the lengths of the past and prediction windows, respectively. The parts colored green are provided as inputs, while the orange and blue parts (with question marks) are outputs. The Flood Manager and Flood Evaluator represent deep learning (DL) models, the former to predict control schedules of controllable hydraulic structures (e.g., gates and pumps) to pre-release water, and the latter to predict the resulting water levels for those control schedules. Loss functions, $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$, penalize the flooding and water wastage beyond pre-specified thresholds, respectively.
The flood mitigation problem is usually treated as an optimization task (Karimanzira 2016), where variables such as control schedules for hydraulic structures are optimized to achieve the desired water levels to effectively mitigate flood risks (Zarei et al. 2021). Random initialization followed by soft-computing techniques such as genetic algorithm (GA) (Leon et al. 2020) is often employed to perform the optimizations. Subsequently, physics-based simulator models such as HEC-RAS and SWMM are used to assess generated control schedules (Sadler et al. 2019; Leon et al. 2014; Yin et al. 2023b). However, such methods are prohibitively slow since they require thousands of time-consuming physics-based simulations (Jafarzadegan et al. 2023).
Over the past decade, machine learning (ML) has become an increasingly powerful tool in various aspects of flood management, including flood prediction (Shi et al. 2023c), flood detection (Tanim et al. 2022), susceptibility assessment (Saha et al. 2021), and post-flood management (Munawar et al. 2019). Despite these advances, ML-based methods have yet to be widely applied to flood mitigation tasks. We address this gap by introducing a fully machine learning-based framework, comprising a generator to generate control schedules and a simulator to assess them. The ML-based framework offers significantly faster response times than compute-intensive physics-based approaches, enabling real-time flood management. Furthermore, the proposed ML-based method allows the simulator to actively guide the generator through gradient descent, thus seamlessly integrating the generator and simulator components and leveraging the differentiability of the learned simulator model (Jyothir et al. 2023). To this end, we propose Forecast Informed Deep Learning Architecture, FIDLAr, for flood management. FIDLAr’s characteristics are summarized as follows:
@@
  • FIDLAr seamlessly combines two DL models in series: Flood Manager and Flood Evaluator. The former model is responsible for generating water pre-release schedules, while the latter model accurately forecasts the resulting water levels. Moreover, with the gradient-based planning, and differentiability of trained Evaluator, it can reinforce the Manager to generate better schedules.
@@
  • FIDLAr is a data-driven approach, learning flood mitigation strategies from historically observed data. Once trained, it offers rapid response capabilities, highlighting the advantages of DL-based models over physics-based models, particularly for real-time flood management.
@@
  • FIDLAr is a model-agnostic framework, where both the Manager and Evaluator could be any type of DL model trained with differentiable loss functions that allow back-propagation.
@@
  • FIDLAr is trained with a customized loss function to balance flood risks and water wastage at the same time, which is a multitask learning method.
2 Related Work
Flood Prediction.
Physics-based mechanic models (e.g., HEC-RAS, SWMM) have been widely used to simulate water levels and flows in river systems (Peker et al. 2024; Rahman and Ali 2024; Gomes Jr et al. 2023; Rivett et al. 2022). However, these models are computationally inefficient and fall short of capturing precise knowledge of study domains (Bentivoglio et al. 2022). Therefore, diverse machine learning (ML) and deep learning (DL) models have been studied as surrogates to simulate water levels and flows. For example, support vector machines (SVMs) have been used to predict the urban flash floods (Yan et al. 2018; Choubin et al. 2019), multivariate regression models were adopted to estimate flood volumes and peak flows (Yang and Chang 2020), random forests and K-nearest neighbors were explored for urban flood inundation mapping (Castro-Gama et al. 2014), gaussian process learning models for fast and accurate flood inundation simulation (Fraehr et al. 2023). Furthermore, deep learning models, such as recurrent neural networks, convolutional neural networks, and Transformers, have been employed for flood inundation (Zhou et al. 2021) and flood prediction (Shi et al. 2023b, c).
Flood Mitigation.
Flood mitigation seeks control schedules of those hydraulic structures to avoid or mitigate flood risks, which requires flood prediction models as simulators to evaluate the control schedules. Researchers have attempted to leverage the genetic algorithm and the pattern search to generate control schedules and physics-based models (e.g., EPA-SWMM5, HEC-RAS) as water simulators (Sadler et al. 2019; Leon et al. 2014, 2020). The Lake Mendocino Operations (LMO) model was developed to simulate operations of Lake Mendocino such as release constraints for flood control and water supply operations (Delaney et al. 2020). However, such methods are computationally intensive since they require thousands of time-consuming simulation trials with physics-based models (Jafarzadegan et al. 2023). Additionally, the genetic algorithm and pattern search techniques are usually heuristic, without positive feedback or guidance to better control those hydraulic structures.
@@ 3 Problem Formulation
Flood management or mitigation aims to manage water levels before extreme weather events. It involves predicting control schedules for hydraulic structures such as gates and pumps within the river system, denoted as $X^{gate,pump}_{t+1:t+k}$, spanning $k$ time points into the future from $t+1$ to $t+k$. This prediction takes as input historical data on all possible variables (see Figure 5), $X$, from the preceding $w$ time points, in conjunction with reliably forecasted covariates (such as rainfall and tide) for the next $k$ time points. Then we could train a deep learning (DL) model, $\mathcal{M}_{\theta}$, with parameters $\theta$:
$\mathcal{M}_{\theta_{M}}:(X^{all}_{t-w+1:t},X^{cov}_{t+1:t+k})\rightarrow X^{gate,pump}_{t+1:t+k},$ (1)
@@ 4 Methodology
4.1 Overview
Intuitively, an ML model can be trained to learn the function $\mathcal{M}_{\theta_{M}}$. However, a key challenge lies in the historical data, which often reflects control schedules that led to flooding or other suboptimal outcomes, making it unsuitable as ground-truth data for traditional supervised learning. To overcome this, we train the model, $\mathcal{M}_{\theta_{M}}$, by leveraging the differentiability of a learned simulator model, $\mathcal{E}_{\theta_{E}}$. First, we pre-train an independent and accurate Flood Evaluator using extensive historical data to model the consequences (e.g., water levels) of various actions (e.g., control schedules). We then frame the control schedule planning in Flood Manager as an optimization problem, seeking actions that minimize undesirable outcomes - floods or water wastage. Both the Evaluator and Manager are implemented as neural networks, with the framework illustrated in Figure 1.
@@ 4.2 Flood Evaluator
Flood Evaluator is tasked with accurately forecasting water levels at designated points of interest within river systems for any control schedule of gates and pumps. The underlying transfer function of the Evaluator is:
$\mathcal{E}_{\theta_{E}}:(X^{all}_{t-w+1:t},X^{cov}_{t+1:t+k},X^{gate,pump}_{t+1:t+k})\rightarrow X^{water}_{t+1:t+k}.$ (2)
@@
The Evaluator is trained independently using large-scale historical data to achieve highly accurate water level predictions for any given set of conditions and control schedules. Therefore, once the Evaluator is well trained, its parameters are frozen while training the Manager, where it plays the role of a trained “referee” - scoring control schedules generated by the Manager by predicting the resulting water levels. It also serves to backpropagate the gradient descent feedback, guiding the Manager to produce more effective control schedules of gates and pumps.
@@
Figure 2: Flood Evaluator. The parts shaded green are used as inputs. Water levels (blue) are the outputs.
@@
Input: recent past data: ${\bf X}^{all}_{t-w+1,t}$
@@
Parameter: $\theta_{E},\theta_{M}$: parameters of Evaluator and Manager; $w,k:$ length of past and prediction windows
1:  // Train Flood Evaluator, $\mathcal{E}_{\theta_{E}}$
@@
4:     $\text{MiniBatch}\leftarrow(\{{\bf X}^{all}_{t-w+1,t},{\bf X}^{cov}_{t+1,t+k}\},{\bf X}^{water}_{t+1,t+k})$
5:     $\hat{{\bf X}}^{water}_{t+1,t+k}\leftarrow\mathcal{E}_{\theta_{E}}({\bf X}^{all}_{t-w+1,t},{\bf X}^{cov}_{t+1,t+k})$
@@
10:  return trained Flood Evaluator, $\mathcal{E}_{\theta_{E}}$
11:  // Train Flood Manager, $\mathcal{M}_{\theta_{M}}$, with frozen $\mathcal{E}_{\theta_{E}}$
@@
14:     $\text{MiniBatch}\leftarrow(\{{\bf X}^{all}_{t-w+1,t},{\bf X}^{rain,tide}_{t+1,t+k}\},{\bf X}^{gate,pump}_{t+1,t+k})$
15:     $\hat{{\bf X}}^{gate,pump}_{t+1,t+k}\leftarrow\mathcal{M}_{\theta_{M}}({\bf X}^{all}_{t-w+1,t},{\bf X}^{rain,tide}_{t+1,t+k})$
16:     $\hat{{\bf X}}^{water}_{t+1,t+k}\leftarrow\mathcal{E}_{\theta_{E}}({\bf X}^{all}_{t-w+1,t},{\bf X}^{rain,tide}_{t+1,t+k},\hat{{\bf X}}^{gate,pump}_{t+1,t+k})$
17:     $\mathcal{L}_{E}=c_{1}\cdot\mathcal{L}_{1}(\hat{{\bf X}}^{water}_{t+1,t+k})+c_{2}\cdot\mathcal{L}_{2}(\hat{{\bf X}}^{water}_{t+1,t+k})$
@@
21:  return trained Flood Manager, $\mathcal{M}_{\theta_{M}}$
4.3 Flood Manager
Flood Manager is to produce control schedules for hydraulic structures (i.e., gates and pumps), taking as inputs reliably predictable future information (rain, tide) and all historical data. Since no ground truth is available, it is trained with the differentiability of the learned Evaluator model. Therefore, we connect the Manager with the Evaluator where the output of Eq. (1) is injected into Eq. (2):
$\mathcal{E}_{\theta_{E}}(X^{all},X^{cov}_{t+1:t+k},\mathcal{M}_{\theta_{M}}(X^{all},X^{cov}_{t+1:t+k}))\rightarrow X^{water}_{t+1:t+k},$ (3)
where $X^{all}=X^{all}_{t-w+1:t}$ and $\theta_{M}$ and $\theta_{E}$ are the parameters of Manager and Evaluator.
The resulting output of water levels can be used to compute the loss in Eq. (6), representing evaluation scores for generated control schedules. Gradient descent (Ruder 2016) can be back-propagated as the feedback to update the parameters of the Manager. The parameter update is presented:
$\theta_{M}:=\theta_{M}-\alpha\cdot\frac{\partial\mathcal{L}}{\partial\theta_{M}},$ (4)
where $\alpha$ is the learning rate and $\frac{\partial\mathcal{L}}{\partial\theta_{M}}$ is the partial derivative of the compound function $\frac{\partial\mathcal{L}}{\partial\theta_{M}}=\frac{\partial\mathcal{L}}{\partial\mathcal{E}}\cdot\frac{\partial\mathcal{E}}{\partial\mathcal{M}}\cdot\frac{\partial\mathcal{M}}{\partial\theta_{M}}$. The training details of FIDLAr are given in Algorithm 1.
Figure 3: Flood Manager. The parts shaded green (historical data) are the inputs, and the parts shaded orange are the outputs. The water levels shaded blue are not predicted.
4.4 Custom Loss Function
Loss functions are critical in steering the learning process. An obvious one is the total time (Figure 4(a)) for which the water levels either exceed the flooding threshold or dip below the water wastage threshold. Another related metric is the extent to which the limits are exceeded to signify the severity of floods or water wastage (Figure 4(b)). The lower threshold for flood management is important in practice, since it prevents water wastage, thereby supporting irrigation, facilitating navigation, and maintaining ecological balance. It also prevents the optimization methods from trivially recommending the depletion of valuable water resources to prevent future flooding. The $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ represent the flooding and water wastage losses, respectively, and the final loss function is a balanced combination as shown in Eq. (6).
where $N$ is the number of water level locations of interest; $k$ is the length of prediction horizon; $X^{flood}$ and $X^{waste}$ represent the thresholds for flooding and water wastage; and the capped version, $\hat{X}^{water}$, is obtained using the Evaluator module. The combined loss function is given by:
$\mathcal{L}_{total}=c_{1}\cdot\mathcal{L}_{1}+c_{2}\cdot\mathcal{L}_{2},$ (6)
where $c_{1}/c_{2}$ dictates the relative importance of $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$.
Figure 4: The two red bars represent a threshold of flooding and a threshold of water wastage. Shown are (a) the time spans when these thresholds are crossed, and (b) the areas between water level curves and threshold bars. Violations of the upper and lower thresholds are captured in $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$.
5 Graph Transformer Network
The Manager and Evaluator modules described so far are model agnostic. We tried many existing architectures for them, as discussed in Section 6.2. We devise the Graph Transformer Network (GTN) architecture by combining graph neural networks (GNNs), attention-based transformer networks, long short-term memory networks (LSTMs), and convolutional neural networks (CNNs). GNN and LSTM modules are combined to learn the spatiotemporal dynamics of water levels, while the Transformer and CNN modules focus on extracting feature representations from the covariates. The attention mechanism (Vaswani et al. 2017) is used to discern interactions between covariates and water levels, as shown in Eq. (7) below. Figure 6 presents the GTN architecture, which is used for both Evaluator and Manager, but with minor changes accordingly of the inputs and outputs (see Figures 2 and 3). We have two GNN layers with 32 and 16 channels, one LSTM layer with 16 units, one CNN with 96 filters, and one Transformer encoder with 3 heads.
$Atte(Q,K,V){}$ $=softmax(\frac{Q^{cov}(K^{water})^{T}}{\sqrt{d}})V^{water}$ (7)
$=softmax(\frac{Q^{water}(K^{cov})^{T}}{\sqrt{d}})V^{cov},$
where $T$ denotes the transpose operation; $water$ and $cov$ represent water levels and covariates; and $d$ is the length of projection embedding where $d=d_{q}=d_{k}=d_{v}=128$.